OpenStudy (anonymous):

Please help me! Compare. Use <, >, or = to complete the statement. A. < B. = C. > I think its A but I'm not sure!

OpenStudy (anonymous):

what is the statement???

OpenStudy (anonymous):

statement..? ??

OpenStudy (midhun.madhu1987):

2^14 and 2^12

OpenStudy (anonymous):

ist=2^14 2nd=2^12 so ist> 2nd